# Statistical Test on Diffusion Model-based Anomaly Detection by Selective Inference

### v3

@@ Results.
5.2 Real Data Experiments
We examined the medical image experiment using the dataset extracted from the dataset in Buda et al. (2019). This dataset includes brain images without tumors as normal images (941 images) and brain images with tumors as abnormal images (939 images). A random selection of 100 images from the normal images was made to estimate the mean and variance. The remaining images were standardized using the estimated mean and variance, and 329 normal images were extracted for training the diffusion model. The model was trained with $T=1000$ and the initial time step of the reverse process was set at $T^{\prime}=300$, with reconstruction performed through 5 step samplings. We set the threshold $\lambda=0.9$ and the kernel size of the averaging filter to $3$. The results of the DMAD-test and naive are shown in the Figure 6. The left column represents the results for normal brain images without tumors, while the right column represents the results for abnormal brain images with tumors. The $p_{\rm selective}$ calculated by the proposed DMAD-test is high for normal images (True Negative) and low for abnormal images (True Positive), indicating that the results are desirable. On the other hand, the $p_{\rm naive}$ obtained by the naive test is low not only for abnormal images but also for normal images (False Positive), indicating the invalidness of the naive test. This result indicates that the DMAD-test detected anomalous regions as statistically significant while avoiding misidentification of normal image as anomaly.
$p_{\rm naive}=0.019$, $p_{\rm selective}=0.643$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.379$ Refer to caption $p_{\rm naive}=0.002$, $p_{\rm selective}=0.290$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.822$ Refer to caption $p_{\rm naive}=0.018$, $p_{\rm selective}=0.373$ Refer to caption $p_{\rm naive}=0.024$, $p_{\rm selective}=0.932$ Refer to caption $p_{\rm naive}=0.002$, $p_{\rm selective}=0.225$
(a) Brain images without tumors
$p_{\rm naive}=0.000$, $p_{\rm selective}=0.001$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.002$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.021$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.028$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.027$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.006$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.016$
(b) Brain images with tumors
@@ 5.2 Real Data Experiments
Figure 6: An example of the results for applying the proposed DMAD-test and the naive test (an invalid test ignoring that the anomalous region was identified by the diffusion model) to brain images.
6 Conclusions
In this study, we proposed a statistical test for anomalous regions in medical images detected by a diffusion model. With the proposed DMAD-test, the false detection rate can be controlled with the significance level because statistical inference is conducted conditional on the fact that the anomalous regions are identified by a diffusion model. We believe this study marks a step toward bridging the gap between generative AI and rigorous statistical inference in medical imaging analysis.
Acknowledgement
Appendix A Accelerated Reverse Processes
Methods for accelerating the reverse process have been proposed in DDPM, DDIM (Song et al., 2022). When taking a strictly increasing subsequence $\tau$ from $\{1,\cdots,T\}$, it is possible to skip the sampling trajectory from $\mathbf{x}_{\tau_{i}}$ to $\mathbf{x}_{\tau_{i-1}}$. In this case, equations (2) and (4) can be rewritten as
where
Appendix B Comparison Methods for Numerical Experiments
We compared our proposed method with the following methods:
  • DMAD-test: The proposed method uses the parametric programming.
  • DMAD-test-oc: The proposed method uses over-conditioning.
  • naive: The naive method. This method uses a conventional $z$-test without any conditioning. The naive $p$-value is calculated as
    $p_{\mathrm{naive}}=\mathbb{P}_{\mathbb{H}_{0}}\left(|Z|>|z_{\mathrm{obs}}|\right).$
    $p_{\mathrm{bonferroni}}=\min(1,2^{n}\cdot p_{\mathrm{naive}}).$
    • Calculate the observed test statistic $z_{\mathrm{obs}}$ by applying the observated image $\bm{X}_{\mathrm{obs}}$ to the diffusion model.
    • The permutation $p$-value is then determined as
      $p_{\mathrm{permutation}}=\frac{1}{B}\sum_{b\in[B]}\bm{1}\{|z^{(b)}|>|z_{\mathrm{obs}}|\},$
      where $\bm{1}\{\cdot\}$ denotes the indicator function.
Appendix C Architecture of the U-Net
Figure 7 shows the architecture of the U-Net used in our experiments. The U-Net has three skip connections, and the Encoder and Decoder blocks.
Figure 7: The architecture of the U-Net

### v4

@@ Appendix F Robustness of the proposed method
Appendix G Experimental settings for the MVTec AD dataset
G.1 Experimental settings of Brain Trumor Segmentation 2023 dataset
We evaluate our method on T2-FLAIR MRI brain scans from the Brain Trumor Segmentation 2023 dataset Karargyris et al. (2023); LaBella et al. (2023). T2-FLAIR MRI which comprises 934 non-skull-stripped 3D scans with dimensions of $240\!\times\!240\!\times\!155$. From these scans, we extracted 2D $240\!\times\!240$ axial slices at axis $95$, resized them to $64\!\times\!64$ pixels, and categorized them based on the anomaly annotations into $532$ normal images (without tumors) and $402$ abnormal images (with tumors). For each scan, we estimated the mean and variance from pixel values excluding both the non-brain regions and tumor regions identified in the ground truth, followed by standardization. We randomly selected $312$ normal images for model training. The model was trained with $T=1000$ and the initial time step of the reverse process was set at $T^{\prime}=300$, with reconstruction performed through 5 step samplings. We set the threshold $\lambda=0.6$ and the kernel size of the averaging filter to $3$. Note that, when testing images of the MRI brain scans, the non-brain regions are not treated as anomalous regions $\mathcal{M}_{\bm{X}}$.
G.2 Experimental settings of Brain Trumor Segmentation 2023 dataset
We evaluate our method on the MVTec AD dataset Bergmann et al. (2019), which consists of $15$ object categories. Each category provides a training set of normal images and a test set containing both normal and abnormal images, with image resolutions ranging from $900\times 900$ to $1024\times 1024$ pixels. For our experiments, we select six categories (bottle, cable, grid) and resize all images to $128\times 128$ pixels. For the type I error rate experiments, we randomly select $50$ normal images from each category, and for the power experiments, we select $50$ abnormal images per category. The diffusion model is trained on the remaining normal images with $T=1000$ total diffusion steps, of which the first $T^{\prime}=300$ steps are used for reconstruction using $4$ sampling steps. We apply an averaging filter with a kernel size of $3$. We set the anomaly threshold $\lambda$ to $1.0$ for bottle and to $1.2$ for cable and grid. To demonstrate in the power experiments, we compute the intersection between the anomalous region detected by the diffusion model and the anomaly annotation. Since the images in the MVTec AD dataset are RGB, we redefine the image data as $\tilde{\bm{X}}\in\mathbb{R}^{hw\times 3}$, where $h$ and $w$ denote the image height and width. We then vectorize each image by
$\bm{X}$ $=\mathrm{vec}(\tilde{\bm{X}})$ (38)
$=(X_{1,1},X_{1,2},X_{1,3},\,X_{2,1},X_{2,2},X_{2,3},\dots,X_{hw,1},X_{hw,2},X_{hw,3})^{\top}\in\mathbb{R}^{n},$
Similarly, we define the reference image $\tilde{\bm{X}}^{\mathrm{ref}}\in\mathbb{R}^{hw\times 3}$ as the average of the training images, and vectorize it in the same way as above.
$\bm{X}^{\mathrm{ref}}$ $=\mathrm{vec}(\tilde{\bm{X}}^{\mathrm{ref}})$ (39)
$=(\tilde{X}_{1,1},\tilde{X}_{1,2},\tilde{X}_{1,3},\,\tilde{X}_{2,1},\tilde{X}_{2,2},\tilde{X}_{2,3},\dots,\tilde{X}^{\mathrm{ref}}_{hw,1},\tilde{X}^{\mathrm{ref}}_{hw,2},\tilde{X}^{\mathrm{ref}}_{hw,3})^{\top}\in\mathbb{R}^{n}$
where $n=3hw$, and accordingly redefine the test statistic in (11) as
$T(\bm{X},\bm{X}^{\mathrm{ref}})=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\sum_{j\in[3]}\tilde{X}_{i,j}-\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\sum_{j\in[3]}\tilde{X}^{\mathrm{ref}}_{i,j}$ (40)
With this setup, the our method can be applied in the same way as in §3.
Appendix H Experimental results for the real-world datasets
In this section, we show the experimental results for the real-world datasets. We applied the proposed proposed to the MRI images and MVTec AD dataset and compared with the naive method.
$p_{\rm naive}:0.031$, $p_{\rm selective}:0.266$ Refer to caption $p_{\rm naive}:0.031$, $p_{\rm selective}:0.331$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.766$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.360$ Refer to caption $p_{\rm naive}:0.031$, $p_{\rm selective}:0.646$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.753$ Refer to caption $p_{\rm naive}:0.011$, $p_{\rm selective}:0.768$
MRI images without tumors
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.026$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.012$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.007$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.003$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.002$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
MRI images with tumors
@@ Appendix H Experimental results for the real-world datasets
Figure 9: An example of the results for applying the proposed proposed and the naive test (an invalid test ignoring that the anomalous region was identified by the diffusion model) to MRI images.
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.907$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.010$
(a) Bottle (Left: Normal, Right: Anomaly)
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.567$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
(b) Cable (Left: Normal, Right: Anomaly)
$p_{\rm naive}:0.028$, $p_{\rm selective}:0.789$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
(c) Grid (Left: Normal, Right: Anomaly)
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.769$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
(d) Transistor (Left: Normal, Right: Anomaly)
Figure 10: An example of applying the proposed method and the naive method to six categories of the MVTec AD dataset (Bottle, Cable, Grid, Transistor). For each category, the left figure shows a normal image and the right figure shows an anomalous image. The proposed $p_{\rm selective}$ remains high for normal samples (True Negatives) and low for anomalous samples (True Positives), demonstrating accurate control of the false detection rate. In contrast, the $p_{\rm naive}$ yields low $p$-values across both normal and anomalous images, indicating an inflated false positive rate and invalidity of the naive method.